OpenStudy (anonymous):

MEDAL and FAN plz help :( Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days. 1) Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. 2) Graph the price function for Stock D. Explain the key features of the graph with complete sentences. 3) A drop in sales has affected Stock D with the function g(x) = -6. Explain to Gordon how Stock D's new price function

OpenStudy (anonymous):

f(x)+g(x) will be created. Graph f(x) + g(x). 4) Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plain. -f(x+2) -f(-x) -f(x) + 3

OpenStudy (anonymous):

A function is a relationship that relates the input and output (sometime the Y and X axis)

OpenStudy (anonymous):

Stock A price f(x) = 25(1.08)^x Stock B price: x ----f(x) 0-----22.00 1-----21.56 2-----21.13 Stock c price: the initial price of this stock is $30, but has been increasing 4% each day.

OpenStudy (anonymous):

I got the first part f(x)=22+1.2 but I'm stuck on the others : /

OpenStudy (anonymous):

@Nnesha do you know how to do this?

Nnesha (nnesha):

where you stuck at ??

Nnesha (nnesha):

i mean which part

OpenStudy (anonymous):

Part B

OpenStudy (anonymous):

well question 2 my answer for question 1 is f(x) = 22 +1.2

Nnesha (nnesha):

they gave you equation and table ???

Nnesha (nnesha):

http://openstudy.com/study#/updates/52bf2eb0e4b0fef2cb9e2264 someone already answered this question

Nnesha (nnesha):

??